OpenStudy (anonymous):

integrate dx/(5+4cosx)

hartnn (hartnn):

did u try standard substitution t= tan (x/2) ??

hartnn (hartnn):

\(\\ \text{ To integrate} \huge \frac{1}{a\sin\:x+b\cos\:x+c}\\ \text{put, t=tan(x/2),then} \large \sin\:x =\frac{2t}{1+t^2} \quad \cos\:x=\frac{1-t^2}{1+t^2} \quad dx=\frac{2}{1+t^2} \\ \)

OpenStudy (anonymous):

thanks well done.

hartnn (hartnn):

hmm....welcome ^_^